# Quantum Multiple Eigenvalue Gaussian filtered Search: an efficient and versatile quantum phase estimation method

### 01 Feb 2024 v1

@@ 2.1 Comparison with previous work
Quantum subspace diagonalization (QSD), quantum Krylov, matrix pencil, and ESPRIT methods, as highlighted in studies such as [CG22, ${}^{+}$20, ${}^{+}$22, MKSCdJ17, ${}^{+}$20, PM19, SY21, SHE20, OTT19, ELN22, ${}^{+}$23], offer an alternative way to solve the eigenvalue estimation problem. These methods estimate eigenvalues by addressing specific projected eigenvalue problems or singular value problems and have proven valuable for estimating ground-state and excited-state energies across various scenarios. Despite classical perturbation theories suggesting potential challenges, such as ill-conditioned projected problems and sensitivity to noise, empirical observations indicate that these quantum methods often outperform pessimistic theoretical predictions. Recently, [SHT22] proposed and investigated the complexity of a quantum phase estimation algorithm based on ESPRIT [RK89, LLF20]. It recovers the frequencies by computing the SVD of the Hankel matrix generated from samples of the signal. Assuming a sufficiently large dominant spectral gap $\Delta_{\rm dom}$ and $p_{\mathrm{tail}}=0$, [SHT22] establishes that ESPRIT can achieve $\epsilon$ precision with $T_{\max}=\widetilde{\cal O}(\mathrm{poly}(|\mathcal{D}|)\epsilon^{-o(1)})$ and $T_{\rm total}=\widetilde{\cal O}(\mathrm{poly}(|\mathcal{D}|)(p_{\min}\epsilon)^{-2})$. Notably, the relationship $T_{\rm total}=\Omega(T_{\max}^{2})$, which arises from the aliasing issue, poses a challenge to the ESPRIT method (as well as other ESPRIT-like methods, such as the observable dynamical mode decomposition (ODMD) method [${}^{+}$23]) in achieving the Heisenberg- limited scaling (see Appendix B for detail). More recently, [LNY23a, Section IV] presented a multilevel ESPRIT technique to overcome the restriction of $p_{\mathrm{tail}}=0$ and reach the Heisenberg limited scaling with short circuit depth (Properties (1), (2), (4)). Specifically, when $\epsilon\ll\Delta_{\text{dom}}$, the authors illustrate that multilevel ESPRIT can achieve $\epsilon$-accuracy with $T_{\max}=\widetilde{\mathcal{O}}(\mathrm{poly}(|\mathcal{D}|)p_{\mathrm{tail}}(p_{\min}\epsilon)^{-1})$ and $T_{\mathrm{total}}=\widetilde{\mathcal{O}}(\mathrm{poly}(|\mathcal{D}|)(p_{\mathrm{tail}}p_{\min}\epsilon)^{-1})$. In addition, despite the apparent importance of the gap assumption in the previous analysis of ESPRIT [SHT22, LNY23a], our numerical experiments in Section 4 suggest that this assumption might be relaxed in real-world applications.
@@
It should be noted that filtering techniques have also been explored in previous studies. For example, in [${}^{+}$23], the authors introduced the Gaussian derivative filter to estimate the ground-state energy (smallest eigenvalues) and achieve a “Constant” depth property (Property (5)). However, it remains unclear how to extend their algorithms to address the MEE problem while satisfying the first four properties (Properties (1)-(4)). More recently,  [${}^{+}$23] proposed an algorithm that employs a filter function based on the derivative Heaviside function. In this approach, the eigenvalues of $H$ can be approximated by identifying the local maxima of the approximate derivative of the cumulative distribution function (CDF). As $T_{\max}$ tends to infinity, their approach aligns with QMEGS in that both methods aim to identify the local maximal point in the sum of delta functions. However, when aiming to identify dominant eigenvalues with finite $T_{\max}$, the performance of their algorithm and its ability to achieve Heisenberg-limited scaling are unclear.
@@
[LNY23a, Theorem III.5] $\operatorname{poly}(|{\cal D}|)$ quantum cost
[LNY23a, Theorem V.1]
MM-QCELS [DL23b]
@@
Table 1: Comparison of the existing theoretical analysis of different methods for multiple eigenvalue estimation. For simplicity, certain properties in this table represent a slightly relaxed version of those introduced in Section 1. Specifically, Allow $p_{\mathrm{tail}}>0$ means that the algorithm can deal with an imperfect initial state. The Heisenberg limit implies that the algorithm can achieve $\epsilon$-accuracy with $T_{\mathrm{total}}=\widetilde{\mathcal{O}}(1/\epsilon)$, assuming that all other parameters remain constant. No gap requirement means that the algorithm can achieve $\epsilon$-accuracy with $T_{\mathrm{total}}=\operatorname{poly}(1/\epsilon)$, where the polynomial and constants are independent of $\Delta,\Delta_{\text{dom}}$. “Short” depth means that, when $\epsilon=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min},|\mathcal{D}|}(\Delta_{\text{dom}})$, the algorithm can achieve $\epsilon$-accuracy with $T_{\max}=\widetilde{\mathcal{O}}_{p_{\min},|\mathcal{D}|}\left(p_{\mathrm{tail}}/\epsilon\right)$ and $T_{\rm total}=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min},|\mathcal{D}|}\left(1/\epsilon\right)$. A more detailed comparison can be found in Section 2.1.
@@ 3 Statement of the main results
    The above theorem guarantees that Properties (1)-(3) are attained as outlined in Section 1. This result shares similarities with [LNY23a, Theorem III.5]. Since a spectral gap is not assumed, $T$ may not be large enough to distinguish between close dominant eigenvalues, and it is possible for two dominant eigenvalues to fall within the same interval $[\boldsymbol{\theta}_{k}-\alpha/T,\boldsymbol{\theta}_{k}+\alpha/T]$. Also, some intervals may not include any dominant eigenvalues since $K\geq|\mathcal{D}|$. But (12) ensures that the dominant eigenvalues are covered by the union set $\cup_{k}[\boldsymbol{\theta}_{k}-\alpha/T,\boldsymbol{\theta}_{k}+\alpha/T]$.
@@
    $G(\boldsymbol{\theta}_{k})\leq p_{\min}+\exp\left(-\frac{\alpha^{2}}{18}\right)+\text{small error}<p_{\min}\exp\left(-\frac{q^{2}}{2}\right)+\text{small error}\leq G(\boldsymbol{\theta}^{\star})\,,$
@@
  • Gapped dominant eigenvalues case: In this case, we assume that $T$ is large enough to allow the Gaussian filter to distinguish dominant eigenvalues. We show that Algorithm 2 attains Heisenberg-limited scaling with $T_{\max}=p_{\mathrm{tail}}/(p_{\mathrm{min}}\epsilon)$. Similar results have been attained by previous algorithms, such as MM-QCELS [DL23b] and [LNY23a, Theorem IV.2], and are referred to as Short depth regime (Property (4)). The results are summarized in the following theorem:
@@
[…]
@@
[…]
@@
  • The case where a gap is known between dominant eigenvalues and the tail: In this case, we assume that $T\gg\Delta^{-1}$. The Gaussian filter thus becomes sufficiently sharp to distinguish between dominant eigenvalues and tail eigenvalues. In this scenario, we demonstrate that Section 2 can achieve Heisenberg-limited scaling, with $T_{\max}\epsilon$ arbitrarily close to zero. Furthermore, Section 2 is capable of achieving $\epsilon$ accuracy with $T_{\max}=\mathcal{O}(\log(1/\epsilon))$ and $T_{\mathrm{total}}=\mathcal{O}(1/\epsilon^{2})$. Since the maximum running time $T_{\max}$ depends only logarithmically on $\epsilon$, we refer to this regime as the “Constant” depth regime (Property (5)). It should be noted that similar results have been achieved previously in the task of ground-state energy estimation algorithms, such as QCELS [DL23a, DDTL23], and [${}^{+}$23]. Here, we extend this result to the case of multiple eigenvalue estimation.
@@
    The algorithm’s ability to achieve a “Constant” depth comes from the rapid decay of the filter function $F(x)$ as $|x|$ increases. A similar concept is employed in QCELS [DL23a, DDTL23], where solving the optimization problem is nearly equivalent to a filtering and searching process using the Gaussian filter function. It is also important to note that the choice of the Gaussian filter function might not be exclusive to achieve the “Constant" depth property. For instance, in the algorithm presented in [${}^{+}$23], the filtering and searching process is applied using the Gaussian derivative function to approximately locate the ’zero point’ of the filtered signal, which corresponds to the ground state energy. Exploring the possibility of extending other filter functions to MEE and achieving similar results would be an interesting avenue for further investigation.
@@ 4 Numerical experiments
In this section, we demonstrate the efficiency of QMEGS numerically by comparing it with previously established quantum phase estimation algorithms: MM-QCELS [DL23b], QPE (textbook version [NC00]), and ESPRIT [SHT22]. A brief summary of these three methods is provided in Appendix B. We share the code on Github (https://github.com/zhiyanding/phase_estimation_methods).
@@
In our test, we construct an initial state |$\ket{\psi}$ with $p_{1}=p_{2}=0.4$. Therefore, we let $\lambda_{1},\lambda_{2}$ be the dominant eigenvalues and set $\mathcal{D}=\{1,2\}$. In addition, the dominant spectral gap is $\Delta_{\text{dom}}=\lambda_{1}-\lambda_{0}$. To test the four algorithms, we set $T=100\times 2^{n}$ with $n=1,2,\cdots,7$. For QMEGS (Section 2), we choose $N=500$, $K=2$, $\alpha=5$, $\sigma=1$, $q=0.05$. For MM-QCELS [DL23b, Algorithm 2], the parameters are set to $K=2,T_{0}=100,N_{0}=10^{3},N_{j\geq 1}=500$ and $\sigma=1$. For ESPRIT (see Appendix B), we set $N=\lfloor T\rfloor$ and $K=2$. For QPE, we sample its distribution $N_{\mathrm{QPE}}=30$ times to estimate $\lambda_{0}$.
We note that the textbook version of QPE [NC00] is designed to estimate the ground-state energy of $H$. It is not straightforward to generalize the procedure (discussed in Appendix B) to estimate multiple eigenvalues. Consequently, in our experiment, we first use QPE to estimate the smallest eigenvalue $\lambda_{1}$ and measure the error accordingly. We then use QMEGS (Section 2), MM-QCELS, and ESPRIT to estimate the two dominant eigenvalues and measure the max-min error:
@@
In our case, $D=\{1,2\},K=2$. While QPE’s error is determined by a single eigenvalue estimation, the error assessment for QMEGS (Section 2), MM-QCELS, and ESPRIT involves the maximum error across two eigenvalues. This testing framework inherently provides QPE with an advantage. Despite this inherent bias, we illustrate that QMEGS (Section 2) has the ability to surpass QPE in performance.
@@ 4.2 Ising model
We observe that the errors of QMEGS, MM-QCELS, and QPE exhibit a proportionality to the inverse of the maximal evolution time ($T_{\max}$). In particular, the constant factor $\delta=T\epsilon$ for QMEGS and MM-QCELS is considerably smaller than that of QPE. As illustrated in Fig. 3, QMEGS and MM-QCELS significantly reduce the maximal evolution time by two orders of magnitude. In contrast to the other three methods, the error of ESPRIT decreases with $T^{-1.5}_{\max}$, enabling it to achieve a smaller error than the other methods when $T_{\max}$ is sufficiently large. However, in terms of the total evolution time $T_{\mathrm{total}}$, ESPRIT incurs a cost that is one order of magnitude higher than QMEGS.
@@ 5 Discussions
  • It is important to emphasize that while ESPRIT-type algorithms [SHT22, ${}^{+}$23] may not achieve Heisenberg-limited scaling and theoretically require gap assumptions, these algorithms still perform effectively in the absence of gaps between dominant eigenvalues and show rapid error decay with $T_{\max}$. In the presence of a gap, numerical results indicate that the decay rate of the error appears to be $\mathcal{O}(T^{-1.5}_{\max})$. To our knowledge this phenomenon was not previously reported in the literature. The theoretical underpinning of these phenomena remains an open question, which makes it an intriguing direction for further exploration into the complexity of ESPRIT-type algorithms.
  • As highlighted in Section 2.1, the Gaussian filtering function is not the only option for addressing the MEE problem. Some other filtering functions, such as the Gaussian derivative filter [${}^{+}$23] and derivative Heaviside function [${}^{+}$23], might also be used to solve the MEE problem. The exploration of alternative filtering functions that offer improved complexity performance and flexibility remains an open problem.

### 26 Sep 2024 v2

@@ 2.1 Comparison with previous work
Quantum subspace diagonalization (QSD), quantum Krylov, matrix pencil, and ESPRIT methods, as highlighted in studies such as [15, 26, 27, 28, 29, 30, 31, 32, 33, 16, 20], offer an alternative way to solve the eigenvalue estimation problem. These methods estimate eigenvalues by addressing specific projected eigenvalue problems or singular value problems and have proven valuable for estimating ground-state and excited-state energies across various scenarios. Despite classical perturbation theories suggesting potential challenges, such as ill-conditioned projected problems and sensitivity to noise, empirical observations indicate that these quantum methods often outperform pessimistic theoretical predictions. Recently, [14] proposed and investigated the complexity of a quantum phase estimation algorithm based on ESPRIT [34, 35]. It recovers the frequencies by computing the SVD of the Hankel matrix generated from samples of the signal. Assuming a sufficiently large dominant spectral gap $\Delta_{\rm dom}$ and $p_{\mathrm{tail}}=0$, [14] establishes that ESPRIT can achieve $\epsilon$ precision with $T_{\max}=\widetilde{\cal O}(\mathrm{poly}(|\mathcal{D}|)\epsilon^{-o(1)})$ and $T_{\rm total}=\widetilde{\cal O}(\mathrm{poly}(|\mathcal{D}|)(p_{\min}\epsilon)^{-2})$. Notably, the relationship $T_{\rm total}=\Omega(T_{\max}^{2})$, which arises from the aliasing issue, poses a challenge to the ESPRIT method (as well as other ESPRIT-like methods, such as the observable dynamical mode decomposition (ODMD) method [20]) in achieving the Heisenberg-limited scaling (see Appendix B for detail). More recently, [19, Section IV] presented a multilevel ESPRIT technique to overcome the restriction of $p_{\mathrm{tail}}=0$ and reach the Heisenberg limited scaling with short circuit depth (Properties (1), (2), (4)). Specifically, when $\epsilon\ll\Delta_{\text{dom}}$, the authors illustrate that multilevel ESPRIT can achieve $\epsilon$-accuracy with $T_{\max}=\widetilde{\mathcal{O}}(\mathrm{poly}(|\mathcal{D}|)p_{\mathrm{tail}}(p_{\min}\epsilon)^{-1})$ and $T_{\mathrm{total}}=\widetilde{\mathcal{O}}(\mathrm{poly}(|\mathcal{D}|)(p_{\mathrm{tail}}p_{\min}\epsilon)^{-1})$. In addition, despite the apparent importance of the gap assumption in the previous analysis of ESPRIT [14, 19], our numerical experiments in Section 4 suggest that this assumption might be relaxed in real-world applications.
@@
It should be noted that filtering techniques have also been explored in previous studies. For example, in [9], the authors introduced the Gaussian derivative filter to estimate the ground-state energy (smallest eigenvalues) and achieve a “Constant” depth property (Property (5)). However, it remains unclear how to extend their algorithms to address the MEE problem while satisfying the first four properties (Properties (1)-(4)). More recently,  [12] proposed an algorithm that employs a filter function based on the derivative Heaviside function. In this approach, the eigenvalues of $H$ can be approximated by identifying the local maxima of the approximate derivative of the cumulative distribution function (CDF). As $T_{\max}$ tends to infinity, their approach aligns with QMEGS in that both methods aim to identify the local maximal point in the sum of delta functions. However, when aiming to identify dominant eigenvalues with finite $T_{\max}$, the performance of their algorithm and its ability to achieve Heisenberg-limited scaling are unclear.
@@
[19, Theorem III.5] $\operatorname{poly}(|{\cal D}|)$ quantum cost
[19, Theorem V.1]
MM-QCELS [18]
@@
Table 1: Comparison of the existing theoretical analysis of different methods for multiple eigenvalue estimation. In the table, a question mark indicates that the method potentially satisfies this property, although there is no existing result that rigorously proves it. For simplicity, certain properties in this table represent a slightly relaxed version of those introduced in Section 1. Specifically, Allow $p_{\mathrm{tail}}>0$ means that the algorithm can deal with an imperfect initial state. The Heisenberg limit implies that the algorithm can achieve $\epsilon$-accuracy with $T_{\mathrm{total}}=\widetilde{\mathcal{O}}(1/\epsilon)$, assuming that all other parameters remain constant. No gap requirement means that the algorithm can achieve $\epsilon$-accuracy with $T_{\mathrm{total}}=\operatorname{poly}(1/\epsilon)$, where the polynomial and constants are independent of $\Delta,\Delta_{\text{dom}}$. “Short” depth means that, when $\epsilon=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min},|\mathcal{D}|}(\Delta_{\text{dom}})$, the algorithm can achieve $\epsilon$-accuracy with $T_{\max}=\widetilde{\mathcal{O}}_{p_{\min},|\mathcal{D}|}\left(p_{\mathrm{tail}}/\epsilon\right)$ and $T_{\rm total}=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min},|\mathcal{D}|}\left(1/\epsilon\right)$. A more detailed comparison can be found in Section 2.1.
@@ 3 Statement of the main results
    The above theorem guarantees that Properties (1)-(3) are attained as outlined in Section 1. This result shares similarities with [19, Theorem III.5]. Since a spectral gap is not assumed, $T$ may not be large enough to distinguish between close dominant eigenvalues, and it is possible for two dominant eigenvalues to fall within the same interval $[\boldsymbol{\theta}_{k}-\alpha/T,\boldsymbol{\theta}_{k}+\alpha/T]$. Also, some intervals may not include any dominant eigenvalues since $K\geq|\mathcal{D}|$. But (12) ensures that the dominant eigenvalues are covered by the union set $\cup_{k}[\boldsymbol{\theta}_{k}-\alpha/T,\boldsymbol{\theta}_{k}+\alpha/T]$.
@@
    $G(\boldsymbol{\theta}_{k})\leq p_{\mathrm{tail}}+\exp\left(-\frac{\alpha^{2}}{18}\right)+\text{small error}<p_{\min}\exp\left(-\frac{q^{2}}{2}\right)+\text{small error}\leq G(\boldsymbol{\theta}^{\star})\,,$
@@
  • Gapped dominant eigenvalues case: In this case, we assume that $T$ is large enough to allow the Gaussian filter to distinguish dominant eigenvalues. We show that Algorithm 2 attains Heisenberg-limited scaling with $T_{\max}=p_{\mathrm{tail}}/(p_{\mathrm{min}}\epsilon)$. Similar results have been attained by previous algorithms, such as MM-QCELS [18] and [19, Theorem IV.2], and are referred to as Short depth regime (Property (4)). The results are summarized in the following theorem:
@@
    $T=\Omega\left(\frac{\alpha}{\Delta_{\text{dom}}}\right),\quad N=\Omega\left(\frac{1}{(\zeta p_{\mathrm{tail}})^{2}}\log\left(\left(\frac{T}{q}+|\mathcal{D}|\right)\frac{1}{\eta}\right)\right)\,.$
@@
    Furthermore, we have $k_{i}\neq k_{j}$ for $i\neq j$. In particular, when $p_{\min}>(1+o(1))p_{\mathrm{tail}}$, for $\epsilon=\widetilde{\mathcal{O}}\left(p_{\mathrm{tail}}\Delta_{\text{dom}}/p_{\min}\right)$, the following quantum cost is enough to achieve $\max_{m\in\mathcal{D}}|\lambda_{m}-\boldsymbol{\theta}_{k_{m}}|\leq\epsilon$:
@@
  • The case where a gap is known between dominant eigenvalues and the tail: In this case, we assume that $T\gg\Delta^{-1}$. The Gaussian filter thus becomes sufficiently sharp to distinguish between dominant eigenvalues and tail eigenvalues. In this scenario, we demonstrate that Section 2 can achieve Heisenberg-limited scaling, with $T_{\max}\epsilon$ arbitrarily close to zero. Furthermore, Section 2 is capable of achieving $\epsilon$ accuracy with $T_{\max}=\mathcal{O}(\log(1/\epsilon))$ and $T_{\mathrm{total}}=\mathcal{O}(1/\epsilon^{2})$. Since the maximum running time $T_{\max}$ depends only logarithmically on $\epsilon$, we refer to this regime as the “Constant” depth regime (Property (5)). It should be noted that similar results have been achieved previously in the task of ground-state energy estimation algorithms, such as QCELS [10, 22], and [9]. Here, we extend this result to the case of multiple eigenvalue estimation.
@@
    The algorithm’s ability to achieve a “Constant” depth comes from the rapid decay of the filter function $F(x)$ as $|x|$ increases. A similar concept is employed in QCELS [10, 22], where solving the optimization problem is nearly equivalent to a filtering and searching process using the Gaussian filter function. It is also important to note that the choice of the Gaussian filter function might not be exclusive to achieve the “Constant" depth property. For instance, in the algorithm presented in [9], the filtering and searching process is applied using the Gaussian derivative function to approximately locate the ’zero point’ of the filtered signal, which corresponds to the ground state energy. Exploring the possibility of extending other filter functions to MEE and achieving similar results would be an interesting avenue for further investigation.
@@ 4 Numerical experiments
In this section, we demonstrate the efficiency of QMEGS numerically by comparing it with previously established quantum phase estimation algorithms: MM-QCELS [18], QPE (textbook version [1]), and ESPRIT [14]. A brief summary of these three methods is provided in Appendix B. We share the code on Github (https://github.com/zhiyanding/phase_estimation_methods).
@@
In our test, we construct an initial state |$\ket{\psi}$ with $p_{1}=p_{2}=0.4$. Therefore, we let $\lambda_{1},\lambda_{2}$ be the dominant eigenvalues and set $\mathcal{D}=\{1,2\}$. In addition, the dominant spectral gap is $\Delta_{\text{dom}}=\lambda_{2}-\lambda_{1}$. To test the four algorithms, we set $T=100\times 2^{n}$ with $n=1,2,\cdots,7$. For QMEGS (Section 2), we choose $N=500$, $K=2$, $\alpha=5$, $\sigma=1$, $q=0.05$. For MM-QCELS [18, Algorithm 2], the parameters are set to $K=2,T_{0}=100,N_{0}=10^{3},N_{j\geq 1}=500$ and $\sigma=1$. For ESPRIT (see Appendix B), we set $N=\lfloor T\rfloor$ and $K=2$. For QPE, we consider the specific version designed for estimating the ground state energy $\lambda_{1}$, which is a variation of the version proposed in  [1, Chapter 5.2]. Specifically, we sample its distribution $N_{\mathrm{QPE}}=\lceil 6/p_{1}\rceil$ times, and take the minimal of the measures as the approximation to $\lambda_{1}$ (see Appendix B). The error QPE is defined as the error for estimating this single eigenvalue.
We then use QMEGS (Section 2), MM-QCELS, and ESPRIT to estimate the two dominant eigenvalues and measure the max-min error:
@@
In our case, $D=\{1,2\},K=2$. This testing criterion gives QPE with an advantage. Nonetheless, we illustrate that the performance of QMEGS, MM-QCELS, and ESPRIT can all to surpass that of QPE.
@@ 4.2 Ising model
We observe that the errors of QMEGS, MM-QCELS, and QPE exhibit a proportionality to the inverse of the maximal evolution time ($T_{\max}$). In particular, the constant factor $\delta=T\epsilon$ for QMEGS and MM-QCELS is considerably smaller than that of QPE. As illustrated in Fig. 3, QMEGS and MM-QCELS significantly reduce the maximal evolution time by two orders of magnitude. In contrast to the other three methods, the error of ESPRIT decreases with $T^{-1.5}_{\max}$ivivivA very recent work [36] rigorously proves this based on a highly nontrivial matrix perturbation argument. Specifically, when $T_{\max}\gg\Delta_{m}$, the authors show that the error of ESPRIT scales as $T_{\max}^{-1.5}$., enabling it to achieve a smaller error than the other methods when $T_{\max}$ is sufficiently large. However, in terms of the total evolution time $T_{\mathrm{total}}$, ESPRIT incurs a cost that is one order of magnitude higher than QMEGS.
@@ 5 Discussions
  • It is important to emphasize that while ESPRIT-type algorithms [14, 20] may not achieve Heisenberg-limited scaling and theoretically require gap assumptions, these algorithms still perform effectively in the absence of gaps between dominant eigenvalues and show rapid error decay with $T_{\max}$. In the presence of a gap, numerical results indicate that the decay rate of the error appears to be $\mathcal{O}(T^{-1.5}_{\max})$. To our knowledge this phenomenon was not previously reported in the literature. The theoretical underpinning of these phenomena remains an open question, which makes it an intriguing direction for further exploration into the complexity of ESPRIT-type algorithms.
  • As highlighted in Section 2.1, the Gaussian filtering function is not the only option for addressing the MEE problem. Some other filtering functions, such as the Gaussian derivative filter [9] and derivative Heaviside function [12], might also be used to solve the MEE problem. The exploration of alternative filtering functions that offer improved complexity performance and flexibility remains an open problem.